assuming the following ratios are constant total assets turnover=2.34% profit margin=6.2% equity multiplier=1.81% payout ratio= 31% what is the sustainable growth rate?
Answer #1

ROE=profit margin*Total asset turnover*Equity multiplier

=(2.34*6.2*1.81)=26.25948%

Retention ratio=1-dividend payout ratio

=1-0.31=0.69

Sustainable Growth Rate=(ROE*Retention ratio)/[1-(ROE*Retention ratio)]
=(0.2625948*0.69)/[1-(0.2625948*0.69)]

which is equal to

=22.13%(Approx).